Two angles are complementary. The sum of the measure of the first angle and half the second angle is 78 deg. Find the measures of the angles.
What is the measure of the smaller angle?
What is the measure of the other angle?

Answer by stanbon(48568)

Two angles are complementary. The sum of the measure of the first angle and half the second angle is 78 deg. Find the measures of the angles.
------------------------
Let the angles be "x" and "90-x"
--------------------------------
EQUATION:
x + (1/2)(90-x) = 78
x + 45 -(x/2) = 78
(1/2)x = 33
x = 66
90-x = 24
--------------------------
What is the measure of the smaller angle?
Ans: 24
What is the measure of the other angle?
Ans: 66
=============
Cheers,
Stan H.